OpenStudy (anonymous):

When solving an exponential equation by graphing (Solve 4^7x+250) on a calculator, I keep getting x and y as zero at the intersection?

OpenStudy (anonymous):

edit: 4^7x=250

OpenStudy (amistre64):

unless you move the graph by subtracting from it.... your never gonna hit the axis on an exponential function. They start from the left at just above zero and grow from there. At x=0, they alway cross at y=1 without modifying the equation.

OpenStudy (amistre64):

your equation only has one intersecting point, (0,251)

OpenStudy (amistre64):

and then theres the edit lol....

OpenStudy (amistre64):

lets log it out.... log4(4^7x) = log4(250) 7x = log4(250) log4(250) x = --------- 7

OpenStudy (amistre64):

your point is going to be: | log4(250) | | -------- , 250 | | 7 |

OpenStudy (anonymous):

Oh I see how that would work. How would you do it on a calculator though? Our teacher said to go into "Y=" and set Y1=4^7x, and Y2=250. But then my graph has only one intersection at 0,0. X cant be zero because if you plug it back into the original equation, you'd get 4^7(0) which is 0=250. not correct....